# 37N35, 65F55, 93A15, 93B15, 93C57

### v1

@@ Abstract
Structured reduced-order modeling is a central component in the computer-aided design of control systems in which cheap-to-evaluate low-dimensional models with physically meaningful internal structures are computed. In this work, we develop a new approach for the structured data-driven surrogate modeling of linear dynamical systems described by second-order time derivatives via balanced truncation model-order reduction. The proposed method is a data-driven reformulation of position-velocity balanced truncation for second-order systems and generalizes the quadrature-based balanced truncation for unstructured first-order systems to the second-order case. The computed surrogates encode a generalized proportional damping structure, and the damping coefficients are inferred solely from data by minimizing a least-squares error over the coefficients. Several numerical examples demonstrate the effectiveness of the proposed method.
@@ 1 Introduction
The mathematical modeling of, e.g., mechanical or electrical structures [45, Ch. 1], typically results in linear dynamical systems described by second-order differential equations of the form
@@
Due to a generally high demand for modeling accuracy, mathematical models such as eq. 1 are typically described by a large number of differential equations, $n\in\mathcal{O}(n^{6})$, making their use in downstream computational tasks such as numerical simulation or optimization very challenging. Simultaneously, in complex applications, explicit state-space models of the form eq. 1 may be either difficult to obtain or wholly unavailable. Instead, the underlying system can only be accessed in the form of data. This motivates data-driven reduced-order modeling, which is the construction of low-dimensional surrogate models solely from system data. In this work, we consider the identification of structured second-order systems eq. 1 directly from input-to-output data. In other words, our goal is to learn a reduced-order model of the form
@@
from evaluations of the transfer function eq. 2, with the reduced-order quantities ${\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{},{\widetilde{\boldsymbol{D}\mkern 2.0mu}\mkern-2.0mu}{},{\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}\in\mathbb{R}^{r\times r}$, ${\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\in\mathbb{R}^{r\times m}$, ${\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{p}},{\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}\in\mathbb{R}^{p\times r}$, ${\widetilde{\boldsymbol{x}}\mkern 2.0mu\mkern-2.0mu}{}\colon[0,\infty)\to\mathbb{R}^{r}$, and $\widetilde{\boldsymbol{y}}\colon[0,\infty)\to\mathbb{R}^{p}$, and a much smaller number of differential equations $r\ll n$.
Second-order systems eq. 1 can always be recast in unstructured (first-order) form by introducing auxiliary variables. Thus, it is in principle always possible to apply classical data-driven modeling techniques to realize unstructured models of the underlying dynamics in eq. 1. However, while it is always possible to write second-order systems in first-order form, the converse is in general not true [26, Sec. III]. Thus, when using unstructured data-driven modeling methods, the underlying internal system structure is typically not retrievable. This is usually an undesired outcome because the internal system structure is what enables the physical reinterpretation of the learned terms, allows for the use of established computational tools, and preserves structure-inherent properties. Further, twice the number of degrees of freedom is required to realize a first-order system that recovers the input-to-output response of the second-order dynamics, and so structured surrogates typically produce more accurate approximations compared to unstructured ones having the same model order. We refer the reader to [45, 37] for comparisons of methods for structured and non-structured surrogate modeling of eq. 1.
@@
  • In Section 3, we introduce the structured data-driven balanced truncation framework that we propose, second-order quadrature-based position-velocity balanced truncation ($\mathsf{soQuadpvBT}$). The framework builds on Theorem 1, which derives the intrusive reduced-order quantities in $\mathsf{sopvBT}$ from frequency-response data. These data are evaluations of the transfer function eq. 2 and its position- and velocity-output subsystems. The proposed method applies to any system that satisfies a generalized proportional damping hypothesis, where $\boldsymbol{D}$ is expressed as a linear combination of $\boldsymbol{M}$ and $\boldsymbol{K}$ with potentially frequency-dependent coefficients $f(s)$ and $g(s)$. Parts of the theoretical results presented in this section were derived in the course of writing the dissertation of the corresponding author [36].
@@
  • Because the damping coefficients $f(s)$ and $g(s)$ are typically not known a priori, in Section 5, we propose a computational procedure to estimate these weights from frequency response data in the special cases of Rayleigh and structural damping. Specifically, Theorem 5 provides the gradients needed to infer suitable damping parameters from data.
@@ 2.1 Generalized proportional damping model
Here, $\boldsymbol{X}\colon\mathbb{C}\to\mathbb{C}^{n}$, $\boldsymbol{U}\colon\mathbb{C}\to\mathbb{C}^{m}$ and $\boldsymbol{Y}\colon\mathbb{C}\to\mathbb{C}^{p}$ are the Laplace transforms of the state, input, and output vectors from eq. 1, respectively, and the internal damping $\boldsymbol{D}\colon\mathbb{C}\to\mathbb{C}^{n\times n}$ is frequency dependent. See, for example, [4, Sec. 2] and [30] for a more detailed discussion of the different types of internal damping that may appear in the modeling of second-order dynamical systems. We note that in the cases where $\boldsymbol{D}$ is constant, the time- and frequency-domain systems $\mathcal{G}$ in eq. 1 and $\mathcal{G}^{\operatorname{s}}$ in eq. 4 are in fact equivalent formulations related to each other via the Laplace transform.
@@ 2.2 Balanced truncation model reduction
where $\boldsymbol{E},\boldsymbol{A}\in\mathbb{R}^{n_{\operatorname{fo}}\times n_{\operatorname{fo}}}$, $\boldsymbol{B}\in\mathbb{R}^{n_{\operatorname{fo}}\times m}$, $\boldsymbol{C}\in\mathbb{R}^{p\times n_{\operatorname{fo}}}$ and $\boldsymbol{q}\colon[0,\infty)\to\mathbb{R}^{n_{\operatorname{fo}}}$. We assume that $\boldsymbol{E}$ is nonsingular, that the realization of the system eq. 8 is controllable and observable, and that the system eq. 8 is asymptotically stable, i.e., that the eigenvalues of the matrix pencil $\lambda\boldsymbol{E}-\boldsymbol{A}$ lie in the open left half of the complex plane. The core idea of $\mathsf{BT}$ is the simultaneous diagonalization of two symmetric positive semi-definite matrices, called system Gramians, and the truncation of states corresponding to small eigenvalues of these Gramians [28, 27]. The system Gramians $\boldsymbol{P},\boldsymbol{E}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Q}\boldsymbol{E}\in\mathbb{R}^{n_{\operatorname{fo}}\times n_{\operatorname{fo}}}$, also known as the controllability and observability Gramians, respectively, are defined via
@@
$\boldsymbol{Q}$ $=\frac{1}{2\pi}\int_{-\infty}^{\infty}(\mathfrak{i}z\boldsymbol{E}-\boldsymbol{A})^{-\mathsf{H}}\boldsymbol{C}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{C}(\mathfrak{i}z\boldsymbol{E}-\boldsymbol{A})^{-1}\operatorname{d}z.$ (9b)
@@ 2.3 Balanced truncation for second-order systems
see, e.g., [33, Prop. 2.1] and [15, Sec. 4.4]. The methods proposed in [26, 34] correspond to balancing selected combinations of the second-order controllability and observability Gramians in eq. 11 with subsequent truncation. In particular, the so-called free-velocity balanced truncation from [26] and the position-velocity balanced truncation ($\mathsf{sopvBT}$) from [34] correspond to the simultaneous diagonalization of $\boldsymbol{P}_{\operatorname{p}},\boldsymbol{Q}_{\operatorname{p}}$ and $\boldsymbol{P}_{\operatorname{p}},\boldsymbol{M}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Q}_{\operatorname{v}}\boldsymbol{M}$, respectively. Algorithmically, balancing and truncation are achieved simultaneously by using an appropriate generalization of the square-root algorithm from the first-order system case [25, 42]. For a more detailed overview of balanced truncation for second-order systems and all the different variations, we refer the reader to [45, 37, 11].
The primary focus of this work lies on the $\mathsf{sopvBT}$ method from [34]. The main steps of the approach are summarized in Algorithm 1. In the case of symmetric second-order systems, i.e., the mass, damping, and stiffness matrices in eq. 1 are $SPD$ with $\boldsymbol{B}_{\operatorname{u}}=\boldsymbol{C}_{\operatorname{p}}^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{C}_{\operatorname{v}}=\boldsymbol{0}$, asymptotic stability is preserved by $\mathsf{sopvBT}$. In fact, in this particular situation, we have that $\boldsymbol{P}_{\operatorname{p}}=\boldsymbol{Q}_{\operatorname{v}}$; see [34, Thm. 3.1]. Note also that Algorithm 1 can be applied if either the position or velocity-output terms in eq. 1 are zero.
@@
7Compute the reduced-order matrices by projection
@@ 3 Data-driven second-order balanced truncation
Along with the integral formulations of the position-reachability and velocity-observability Gramians $\boldsymbol{P}_{\operatorname{p}}\in\mathbb{R}^{n\times n}$ and $\boldsymbol{M}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Q}_{\operatorname{v}}\boldsymbol{M}\in\mathbb{R}^{n\times n}$ in eq. 12, this observation suggests a natural extension of $\mathsf{foQuadBT}$ for the second-order system eq. 1. In [21], the authors arrive at a fully data-driven reformulation of the first-order $\mathsf{BT}$ method by replacing the exact square-root factors of $\boldsymbol{R}$ and $\boldsymbol{L}$ with low-rank factors derived from numerical quadrature rules applied to the integral representations of the Gramians eq. 9. The quadrature-based factors are never explicitly formed and the method only makes reference to transfer function evaluations at the underlying quadrature nodes. We use this idea as a blueprint to develop a data-driven reformulation of $\mathsf{sopvBT}$. Specifically, we use implicit numerical quadrature rules to derive low-rank approximations of the exact factors $\boldsymbol{R}_{\operatorname{p}}$ and $\boldsymbol{L}_{\operatorname{v}}$, and ultimately show how to realize the quadrature-based approximations of the intrusive reduced-order quantities in eq. 14 non-intrusively using only transfer function data.
@@
Replacing the exact Cholesky factors $\boldsymbol{R}_{\operatorname{p}}$ and $\boldsymbol{L}_{\operatorname{v}}$ in Algorithm 1 with the quadrature-based factors ${\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}$ and ${\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}$ from eqs. 16 and 17 in forming the matrices eq. 14 already yields a low-rank implementation of $\mathsf{sopvBT}$. As we show next in Theorem 1, the significant implication of this low-rank formulation is that the quadrature-based approximations to the intrusive reduced-order quantities in eq. 14 can be computed non-intrusively from transfer function samples and evaluations of the damping coefficients $f(s)$ and $g(s)$ in eq. 5.
@@
$\boldsymbol{X}_{\boldsymbol{j},\boldsymbol{k}}\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\boldsymbol{X}_{(k-1)p+1:kp,(j-1)m+1:jm}\in\mathbb{C}^{K\times J},$
@@
To simplify the proof of the main result Theorem 1, we first show the following lemma from [17].
Lemma 1.
For matrices $\boldsymbol{X},\boldsymbol{Y}\in\mathbb{C}^{n\times n}$, the following two identities
@@
hold for any $s,z\in\mathbb{C}$, with $s\neq z$ and such that $s\boldsymbol{X}+\boldsymbol{Y}$ and $z\boldsymbol{X}+\boldsymbol{Y}$ are nonsingular.
Proof.
For any $s,z\in\mathbb{C}$ for which $s\boldsymbol{X}+\boldsymbol{Y}$ and $z\boldsymbol{X}+\boldsymbol{Y}$ are nonsingular, we observe that
$(s-z)(s\boldsymbol{X}+\boldsymbol{Y})^{-1}\boldsymbol{X}(z\boldsymbol{X}+\boldsymbol{Y})^{-1}$
$\leavevmode\nobreak\ =(s\boldsymbol{X}+\boldsymbol{Y})^{-1}\big{(}(s\boldsymbol{X}+\boldsymbol{Y})-(z\boldsymbol{X}+\boldsymbol{Y})\big{)}(z\boldsymbol{X}+\boldsymbol{Y})^{-1}$
$\leavevmode\nobreak\ =(z\boldsymbol{X}+\boldsymbol{Y})^{-1}-(s\boldsymbol{X}+\boldsymbol{Y})^{-1}$
holds, thus proving eq. 19a. Similarly, we observe that
$(z-s)(s\boldsymbol{X}+\boldsymbol{Y})^{-1}\boldsymbol{Y}(z\boldsymbol{X}+\boldsymbol{Y})^{-1}$
$\leavevmode\nobreak\ =(s\boldsymbol{X}+\boldsymbol{Y})^{-1}\big{(}(sz\boldsymbol{X}+z\boldsymbol{Y})-(sz\boldsymbol{X}+z\boldsymbol{Y})\big{)}(z\boldsymbol{X}+\boldsymbol{Y})^{-1}$
$\leavevmode\nobreak\ =z(z\boldsymbol{X}+\boldsymbol{Y})^{-1}-s(s\boldsymbol{X}+\boldsymbol{Y})^{-1},$
holds, which proves eq. 19b. ∎
@@
where $f(s)$ and $g(s)$ are the damping coefficient functions from eq. 5. Suppose the left and right quadrature nodes $\mathfrak{i}\vartheta_{1},\ldots,\mathfrak{i}\vartheta_{K}$ and $\mathfrak{i}\zeta_{1},\ldots,\mathfrak{i}\zeta_{J}$ in eqs. 16 and 17 are such that
@@
$\mathbb{M}$ =$\stackrel{{\scriptstyle\operatorname{def}}}{{=}}{\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}^{\mathsf{H}}\boldsymbol{M}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}\in\mathbb{C}^{pK\times mJ},\qquad$ $\mathbb{K}$ =$\stackrel{{\scriptstyle\operatorname{def}}}{{=}}{\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}^{\mathsf{H}}\boldsymbol{K}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}\in\mathbb{C}^{pK\times mJ},$ (21a)
$\mathbb{B}_{\operatorname{u}}$ =$\stackrel{{\scriptstyle\operatorname{def}}}{{=}}{\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}^{\mathsf{H}}\boldsymbol{B}_{\operatorname{u}}\in\mathbb{C}^{pK\times m},\qquad$ $\mathbb{C}_{\operatorname{p}}$ =$\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\boldsymbol{C}_{\operatorname{p}}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}\in\mathbb{C}^{p\times mJ},$ (21b)
$\mathbb{C}_{\operatorname{v}}$ =$\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\boldsymbol{C}_{\operatorname{v}}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}\in\mathbb{C}^{p\times mJ}.$ (21c)
@@
where $\boldsymbol{e}_{i}\in\mathbb{R}^{n}$ is the $i$-th canonical basis vector with $1$ at the $i$-th position and 00 everywhere else. Note that $\boldsymbol{I}_{i,\ell}$ contains a subset of the columns of the $n\times n$ identity and has the effect of extracting the columns $(i-1)\ell+1$ through $i\ell$ of a matrix by right multiplication. The formulae for $\mathbb{B}_{\operatorname{u}}$, $\mathbb{C}_{\operatorname{p}}$ and $\mathbb{C}_{\operatorname{v}}$ in eq. 22c are a direct consequence of their construction in eq. 21 and the definitions of the quadrature-based factors ${\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}$ and ${\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}$. We easily observe that
[…]

### v2

@@ Abstract
Structured reduced-order modeling is a central component in the computer-aided design of control systems in which cheap-to-evaluate low-dimensional models with physically meaningful internal structures are computed. In this work, we develop a new approach for the structured data-driven surrogate modeling of linear dynamical systems described by second-order time derivatives via balanced truncation model-order reduction. The proposed method is a data-driven reformulation of position-velocity balanced truncation for second-order systems and generalizes the quadrature-based balanced truncation for unstructured first-order systems to the second-order case. The computed surrogates encode a generalized proportional damping structure, and we propose a computational procedure for inferring the damping coefficients from data by minimizing a least-squares error over the coefficients. Several numerical examples demonstrate the effectiveness of the proposed method.
@@ 1 Introduction
The mathematical modeling of, e.g., mechanical or electrical structures [42, Ch. 1], typically results in linear dynamical systems described by second-order differential equations of the form
@@
Due to a generally high demand for modeling accuracy, mathematical models such as eq. 1 are typically described by a large number of differential equations, $n\sim 10^{6}$, making their use in downstream computational tasks such as numerical simulation or optimization very challenging. Simultaneously, in complex applications, explicit state-space models of the form eq. 1 may be either difficult to obtain or wholly unavailable. Instead, the underlying system can only be accessed in the form of data. This motivates data-driven reduced-order modeling, which is the construction of low-dimensional surrogate models solely from system data. In this work, we consider the problem of learning low-dimensional representations of structured second-order systems as in eq. 1 directly from input-to-output data. In other words, our goal is to learn another, so-called reduced-order model of the form
@@
from evaluations of the transfer function eq. 2 with the reduced-order quantities ${\widetilde{\boldsymbol{M}\mkern 2.0mu}\mkern-2.0mu}{},{\widetilde{\boldsymbol{D}\mkern 2.0mu}\mkern-2.0mu}{},{\widetilde{\boldsymbol{K}\mkern 2.0mu}\mkern-2.0mu}{}\in\mathbb{R}^{r\times r}$, ${\widetilde{\boldsymbol{B}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{u}}\in\mathbb{R}^{r\times m}$, ${\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{p}},{\widetilde{\boldsymbol{C}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}\in\mathbb{R}^{p\times r}$, ${\widetilde{\boldsymbol{x}}\mkern 2.0mu\mkern-2.0mu}{}\colon[0,\infty)\to\mathbb{R}^{r}$, and $\widetilde{\boldsymbol{y}}\colon[0,\infty)\to\mathbb{R}^{p}$, and having a much smaller number of differential equations $r\ll n$.
Second-order systems eq. 1 can always be recast in unstructured (first-order) form by introducing auxiliary variables. Thus, it is in principle always possible to apply classical data-driven modeling techniques to realize unstructured models of the underlying dynamics in eq. 1. However, while it is always possible to write second-order systems in first-order form, the converse is in general not true [24, Sec. III]. Thus, when using unstructured data-driven modeling methods, the underlying internal system structure is typically not retrievable. This is usually an undesired outcome because the internal system structure is what enables the physical reinterpretation of the learned terms, allows for the use of established computational tools, and preserves structure-inherent properties. Further, twice the number of degrees of freedom is required to realize a first-order system that recovers the input-to-output response of the second-order dynamics, and so structured surrogates typically produce more accurate approximations compared to unstructured ones having the same model order. We refer the reader to [42, 34] for comparisons of methods for structured and non-structured surrogate modeling of eq. 1.
@@
  • In Section 3, we introduce the structured data-driven balanced truncation framework that we propose, second-order quadrature-based position-velocity balanced truncation ($\mathsf{soQuadpvBT}$). The framework builds on Theorem 1, which derives the intrusive reduced-order quantities in $\mathsf{sopvBT}$ from frequency-response data. These data are evaluations of the transfer function eq. 2 and its position- and velocity-output subsystems. The proposed method applies to any system that satisfies a generalized proportional damping hypothesis, where $\boldsymbol{D}$ is expressed as a linear combination of $\boldsymbol{M}$ and $\boldsymbol{K}$ with potentially frequency-dependent coefficients $f$ and $g$.
@@
  • Because the damping coefficients $f$ and $g$ are typically not known a priori, in Section 5, we propose a computational procedure to recover these weights from frequency response data in the special cases of Rayleigh and structural damping. Specifically, Theorem 5 provides the gradients needed to recover the underlying damping parameters from data.
@@
Parts of the theoretical results presented in Section 3 were derived in the course of writing the dissertation of the corresponding author [33] and published therein. In [41], the authors independently derive similar results to those contained in [33] for the case of second-order velocity balanced truncation. In contrast to [41], our work considers position-velocity balanced truncation and is more general in the sense that it is presented for a broader class of dynamical systems, namely multiple-input/multiple-output second-order systems with position as well as velocity outputs and with the generalized proportional damping model eq. 5. Additionally, we note that most of the considerations in Sections 4 and 5 have not been discussed for the velocity balancing approach in [41], namely the stability preservation and the computational procedure for recovering unknown damping parameters.
@@ 2.1 Generalized proportional damping model
Here, $\boldsymbol{X}\colon\mathbb{C}\to\mathbb{C}^{n}$, $\boldsymbol{U}\colon\mathbb{C}\to\mathbb{C}^{m}$ and $\boldsymbol{Y}\colon\mathbb{C}\to\mathbb{C}^{p}$ are the Laplace transforms of the state, input, and output vectors from eq. 1, respectively, and the internal damping $\boldsymbol{D}\colon\mathbb{C}\to\mathbb{C}^{n\times n}$ is frequency dependent. See, for example, [3, Sec. 2] and [27] for a more detailed discussion of the different types of internal damping that may appear in the modeling of second-order dynamical systems. We note that in the cases where $\boldsymbol{D}$ is constant, the time- and frequency-domain systems $\mathcal{G}$ in eq. 1 and $\mathcal{G}^{\operatorname{s}}$ in eq. 4 are in fact equivalent formulations related to each other via the Laplace transform.
@@ 2.2 Balanced truncation model reduction
where $\boldsymbol{E},\boldsymbol{A}\in\mathbb{R}^{n_{\operatorname{fo}}\times n_{\operatorname{fo}}}$, $\boldsymbol{B}\in\mathbb{R}^{n_{\operatorname{fo}}\times m}$, $\boldsymbol{C}\in\mathbb{R}^{p\times n_{\operatorname{fo}}}$ and $\boldsymbol{q}\colon[0,\infty)\to\mathbb{R}^{n_{\operatorname{fo}}}$. We assume that $\boldsymbol{E}$ is nonsingular, that the realization of the system eq. 8 is controllable and observable, and that the system eq. 8 is asymptotically stable, i.e., that the eigenvalues of the matrix pencil $\lambda\boldsymbol{E}-\boldsymbol{A}$ lie in the open left half of the complex plane. The core idea of $\mathsf{BT}$ is the simultaneous diagonalization of two symmetric positive semi-definite matrices, called system Gramians, and the truncation of states corresponding to small eigenvalues of these Gramians [26, 25]. The system Gramians $\boldsymbol{P},\boldsymbol{E}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Q}\boldsymbol{E}\in\mathbb{R}^{n_{\operatorname{fo}}\times n_{\operatorname{fo}}}$, also called the controllability and observability Gramians, respectively, are defined via
@@
$\boldsymbol{Q}$ $=\frac{1}{2\pi}\int\limits_{-\infty}^{\infty}(\mathfrak{i}z\boldsymbol{E}-\boldsymbol{A})^{-\mathsf{H}}\boldsymbol{C}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{C}(\mathfrak{i}z\boldsymbol{E}-\boldsymbol{A})^{-1}\operatorname{d}z.$ (9b)
@@ 2.3 Balanced truncation for second-order systems
see, e.g., [30, Prop. 2.1] and [13, Sec. 4.4]. The methods proposed in [24, 31] correspond to balancing selected combinations of the second-order controllability and observability Gramians in eq. 11 with subsequent truncation. In particular, the so-called free-velocity balanced truncation from [24] and the position-velocity balanced truncation ($\mathsf{sopvBT}$) from [31] correspond to the simultaneous diagonalization of $\boldsymbol{P}_{\operatorname{p}},\boldsymbol{Q}_{\operatorname{p}}$ and $\boldsymbol{P}_{\operatorname{p}},\boldsymbol{M}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Q}_{\operatorname{v}}\boldsymbol{M}$, respectively. Algorithmically, balancing and truncation are achieved simultaneously by using an appropriate generalization of the square-root algorithm from the first-order system case [23, 39]. For a more detailed overview of balanced truncation for second-order systems and all the different variations, we refer the reader to [42, 34, 9].
The primary focus of this work lies on the $\mathsf{sopvBT}$ method from [31]. The main steps of the approach are summarized in Algorithm 1. In the case of symmetric second-order systems, i.e., the mass, damping, and stiffness matrices in eq. 1 are $SPD$ with $\boldsymbol{B}_{\operatorname{u}}=\boldsymbol{C}_{\operatorname{p}}^{\mkern-1.5mu\mathsf{T}}$ and $\boldsymbol{C}_{\operatorname{v}}=\boldsymbol{0}$, asymptotic stability is preserved by $\mathsf{sopvBT}$. In fact, in this particular situation, we have that $\boldsymbol{P}_{\operatorname{p}}=\boldsymbol{Q}_{\operatorname{v}}$; see [31, Thm. 3.1]. Note also that Algorithm 1 can be applied if either the position or velocity-output terms in eq. 1 are zero.
@@
7Using $\boldsymbol{W}=\boldsymbol{L}_{\operatorname{v}}\boldsymbol{U}_{1}\boldsymbol{S}_{1}^{-1/2}$ and $\boldsymbol{T}=\boldsymbol{R}_{\operatorname{p}}\boldsymbol{V}_{1}\boldsymbol{S}_{1}^{-1/2}$ compute the reduced-order matrices by projection
@@ 3 Data-driven second-order balanced truncation
Along with the integral formulations of the position-reachability and velocity-observability Gramians $\boldsymbol{P}_{\operatorname{p}}\in\mathbb{R}^{n\times n}$ and $\boldsymbol{M}^{\mkern-1.5mu\mathsf{T}}\boldsymbol{Q}_{\operatorname{v}}\boldsymbol{M}\in\mathbb{R}^{n\times n}$ in eq. 12, this observation suggests a natural extension of $\mathsf{foQuadBT}$ for the second-order system eq. 1. In [19], the authors arrive at a fully data-driven reformulation of the first-order $\mathsf{BT}$ method by replacing the exact square-root factors of $\boldsymbol{R}$ and $\boldsymbol{L}$ with low-rank factors derived from numerical quadrature rules applied to the integral representations of the Gramians eq. 9. The quadrature-based factors are never explicitly formed, and the method only makes reference to transfer function evaluations at the underlying quadrature nodes. We use this idea as a blueprint to develop a data-driven reformulation of $\mathsf{sopvBT}$. Specifically, we use implicit numerical quadrature rules to derive low-rank approximations of the exact factors $\boldsymbol{R}_{\operatorname{p}}$ and $\boldsymbol{L}_{\operatorname{v}}$, and ultimately show how to realize the quadrature-based approximations of the intrusive reduced-order quantities in eq. 14 non-intrusively using only transfer function data.
@@
Replacing the exact Cholesky factors $\boldsymbol{R}_{\operatorname{p}}$ and $\boldsymbol{L}_{\operatorname{v}}$ in Algorithm 1 with the quadrature-based factors ${\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}$ and ${\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}$ from eqs. 16 and 17 in forming the matrices eq. 14 already yields a low-rank implementation of $\mathsf{sopvBT}$. As we show next in Theorem 1, the significant implication of this low-rank formulation is that the quadrature-based approximations to the intrusive reduced-order quantities in eq. 14 can be computed non-intrusively from transfer function samples and evaluations of the damping coefficients $f$ and $g$ in eq. 5.
@@
$\boldsymbol{X}_{\boldsymbol{k},\boldsymbol{j}}\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\boldsymbol{X}_{(k-1)p+1:kp,(j-1)m+1:jm}\in\mathbb{C}^{K\times J},$
@@
To simplify the proof of the main result Theorem 1, we recall the well-known resolvent identities
@@
which hold for any matrices $\boldsymbol{X},\boldsymbol{Y}\in\mathbb{C}^{n\times n}$ and $s,z\in\mathbb{C}$ with $s\neq z$ and such that $s\boldsymbol{X}+\boldsymbol{Y}$ and $z\boldsymbol{X}+\boldsymbol{Y}$ are nonsingular; see, e.g., [15].
@@
where $f$ and $g$ are the damping coefficient functions from eq. 5. Suppose the left and right quadrature nodes $\mathfrak{i}\vartheta_{1},\ldots,\mathfrak{i}\vartheta_{K}$ and $\mathfrak{i}\zeta_{1},\ldots,\mathfrak{i}\zeta_{J}$ in eqs. 16 and 17 are such that
@@
$\mathbb{M}$ $\stackrel{{\scriptstyle\operatorname{def}}}{{=}}{\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}^{\mathsf{H}}\boldsymbol{M}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}\in\mathbb{C}^{pK\times mJ},\qquad$ $\mathbb{K}$ $\stackrel{{\scriptstyle\operatorname{def}}}{{=}}{\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}^{\mathsf{H}}\boldsymbol{K}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}\in\mathbb{C}^{pK\times mJ},$ (21a)
$\mathbb{B}_{\operatorname{u}}$ $\stackrel{{\scriptstyle\operatorname{def}}}{{=}}{\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}^{\mathsf{H}}\boldsymbol{B}_{\operatorname{u}}\in\mathbb{C}^{pK\times m},\qquad$ $\mathbb{C}_{\operatorname{p}}$ $\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\boldsymbol{C}_{\operatorname{p}}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}\in\mathbb{C}^{p\times mJ},$ (21b)
$\mathbb{C}_{\operatorname{v}}$ $\stackrel{{\scriptstyle\operatorname{def}}}{{=}}\boldsymbol{C}_{\operatorname{v}}{\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}\in\mathbb{C}^{p\times mJ}.$ (21c)
@@
where $\boldsymbol{e}_{i}\in\mathbb{R}^{n}$ is the $i$-th canonical basis vector with $1$ at the $i$-th position and 0 everywhere else. Note that $\boldsymbol{I}_{i,\ell}$ contains a subset of the columns of the $n\times n$ identity and has the effect of extracting the columns $(i-1)\ell+1$ through $i\ell$ of a matrix by right multiplication. The formulae for $\mathbb{B}_{\operatorname{u}}$, $\mathbb{C}_{\operatorname{p}}$ and $\mathbb{C}_{\operatorname{v}}$ in eq. 22c are a direct consequence of their construction in eq. 21 and the definitions of the quadrature-based factors ${\check{\boldsymbol{R}\mkern 3.0mu}\mkern-3.0mu}{}_{\operatorname{p}}$ and ${\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}$. We easily observe that
$\left(\mathbb{B}_{\operatorname{u}}\right)_{\boldsymbol{k},:}=\boldsymbol{I}_{k,p}^{\mkern-1.5mu\mathsf{T}}\mathbb{B}_{\operatorname{u}}=\boldsymbol{I}_{k,p}^{\mkern-1.5mu\mathsf{T}}{\check{\boldsymbol{L}\mkern 2.0mu}\mkern-2.0mu}{}_{\operatorname{v}}^{\mathsf{H}}\boldsymbol{B}_{\operatorname{u}}=\varphi_{k}\left(\boldsymbol{C}_{\operatorname{p}}+\mathfrak{i}\vartheta_{k}\boldsymbol{C}_{\operatorname{v}}\right)\boldsymbol{\varphi}(\mathfrak{i}\vartheta_{k})^{-1}\boldsymbol{B}_{\operatorname{u}}=\varphi_{k}\boldsymbol{G}(\mathfrak{i}\vartheta_{k})$